please explain all steps7. Suppose lightning strikes at an average of 1.4 strikes per minute during a particular storm. You play the following game:
Answer #1

here for exponential distribution parameter \lambda =1.4

therefore P(next strike occur within next minute)=P(X<1)=1-e-\lambdax =1-e-1.4 =0.7534

P(next strike occur between 1 and 2 minute)=P(!<X<2)=(1-e-1.4*2)-(1-e-1.4*1)=0.1858

P(strike after 2 minutes)= e-1.4*2 =0.0608

therefore expected value =fair price to be charges =3*0.7534+5*0.1858+1*0.0608=$ 3.25